OpenStudy (anonymous):

Find the solution of the equation 4sin^2(theta) - 1 = 0 in the interval [0, pi/2].

OpenStudy (anonymous):

\[4\sin^2(x)=1\] \[\sin^2(x)=\frac{1}{4}\] \[\sin(x)=\pm\sqrt{\frac{1}{4}}\] \[\sin(x)=\frac{\sqrt{2}}{2}, \sin(x)=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

so \[x=\frac{\pi}{4}, \frac{3\pi}{4},\frac{5\pi}{4},\frac{7\pi}{4}\]

OpenStudy (anonymous):

In my homework, all of the denominators are set at six in the four answer choices. Is something missing?